a man is 3 years older than his wife and four time as old as his son if the son became 15 years old after 3 years what is present age of the wife ?

Answer by ankor@dixie-net.com(22740)

a man is 3 years older than his wife
m = (w+3)
:
and four times as old as his son
m = 4s
:
If the son became 15 years old after 3 years what is present age of the wife?
:
Son's present age, 15 - 3 = 12 yrs old
:
Then
m = 4(12)
m = 48 yrs
and
w = 48 - 3
w = 45 is the wife